# LoRA vs Full Fine-tuning: An Illusion of Equivalence

### v1

@@ Appendix I Case Study: Setting alpha=8 instead of alpha=2r
Figure 14: Number of intruder dimensions in RoBERTa models fine-tuned on 6 different tasks. Here, we set $k=10$. We use the same conditions as in Fig. 5(a) but instead now set $\alpha=8$ instead of $\alpha=2r$.
@@
Figure 15: Impact of the number of singular vectors in the fine-tuned matrix we examine, $k$, on the number of intruder dimensions for RoBERTa models fine-tuned on 6 different tasks. Here, we set $\epsilon=0.5$. We use the same conditions as in Fig. 11(a) but instead now set $\alpha=8$ instead of $\alpha=2r$.
@@
Figure 16: For $\alpha=8$. RoBERTa’s performance on its pre-training data distribution after fine-tuning on a particular task. We measure pseudo loss as described by Salazar et al. (2020). We compare these results to when $\alpha=2r$ (Fig. 9).
@@
Figure 18: Effective rank of RoBERTa on MNLI when $\alpha=8$. We compare these results to when $\alpha=2r$ (Fig. 13). We show for all weight types. For a specific weight type, the graph on the left shows the effective rank of all models, and the right shows the effective rank of the LoRA models only.

### v2

@@ Appendix I Impact of Matrix Percentage on Number of Intruder Dimensions
(a) Number of intruder dimensions in RoBERTa models fine-tuned on 6 different tasks. Here, we set $k=10$. We use the same conditions as in Fig. 4(d) but instead now set $\alpha=8$ instead of $\alpha=2r$.
@@
(b) Impact of the number of singular vectors in the fine-tuned matrix we examine, $k$, on the number of intruder dimensions for RoBERTa models fine-tuned on 6 different tasks. Here, we set $\epsilon=0.5$. We use the same conditions as in Fig. 15(a) but instead now set $\alpha=8$ instead of $\alpha=2r$.
Figure 19: We find that when $\alpha=8$ instead of $\alpha=2r$, our models have more intruder dimensions. (Top) Replication of Fig. 4(d) with $\alpha=8$ instead of $\alpha=2r$. (Bottom) Replication of Fig. 15(a) with $\alpha=8$ instead of $\alpha=2r$.
@@
Figure 20: For $\alpha=8$. RoBERTa’s performance on its pre-training data distribution after fine-tuning on a particular task. We measure pseudo loss as described by Salazar et al. [2020]. We compare these results to when $\alpha=2r$ (Fig. 6).
@@
Appendix N Impact of Random Seeds
To ensure that random seed does not play a role in the number of intruder dimensions we observe in a model, we sample 5 different seeds and fine-tune RoBERTa-base on MNLI using the same methodology as in Fig. 4(d). We find that the initialization has a negligible role on the number of intruder dimensions. This shows that our findings are not dependent on the random initialization of the LoRA modules.
Figure 22: Impact of Random Seeds on intruder dimensions. We fine-tune RoBERTa-base across 5 random seeds and use our same methodology as in Fig. 4(d). We find that the initialization has a negligible role on the number of intruder dimensions. This shows that our findings are not dependent on the random initialization of the LoRA modules.
Appendix O LoRA Variants
We focus significantly on the standard LoRA method proposed by Hu et al. [2021] in order to study it in depth. However, many variants of LoRA have been proposed recently. AdaLoRA [Zhang et al., 2023] adaptively allocates LoRA rank to different modules in order to ensure optimal allocation of trainable parameters to certain modules. LoRA+ [Hayou et al., 2024] sets different learning rates for the A and B modules in LoRA. PiSSA [Meng et al., 2024] initializes the A and B modules with the top ranking singular vectors of the pre-trained weights. VeRA [Kopiczko et al., 2024] models LoRA as the product of two random matrices with trainable parameters doing elementwise operations on the resulting vectors. These variants may have important impacts on the presence of intruder dimensions. For example, PiSSA initializes with the singular vectors and therefore may have an easier time changing them, possibly leading to more intruder dimensions. In contrast, LoRA+ in effect lowers the learning rate, which we found to be important to introducing intruder dimensions, and may therefore reduce the number of intruder dimensions.
In order to examine if intruder dimensions are still relevant for these methods, we rerun our MNLI fine-tuning experiment with RoBERTa with each of these methods with default hyperparameters that they provide. These results are supplied in Fig. 23. Interestingly, we see that all the variations we examine have intruder dimensions. Some interesting observations include: LoRA+ and LoRA $r=1$ appear to have nearly identical curves, suggesting they have very ismilar intruder dimension characteristics. We again see that with higher ranks ($r=64$) these LoRA variants tend to have very few intruder dimensions. However, it does appear that methods that explicitly modify the singular vectors, like PiSSA, have many intruder dimensions. This makes sense since they are explicitly constructed to modify the singular vectors on the pre-trained model. These findings emphasize that intruder dimensions are not just an observed phenomenom in normal LoRA and suggests to future work the examination of LoRA variants.
Figure 23: Measuring LoRA variants for intruder dimensions. $k=10$. We compare variants of LoRA to normal LoRA (blue) and full fine-tuning (black). We find that the LoRA variants we examine still have intruder dimensions and shows that our findings are not just exclusive to normal LoRA.